# From Misleading Queries to Accurate Answers: A Three-Stage Fine-Tuning Method for LLMs

### v1

@@ Abstract
Large language models (LLMs) exhibit excellent performance in natural language processing (NLP), but remain highly sensitive to the quality of input queries, especially when these queries contain misleading or inaccurate information. Existing methods focus on correcting the output, but they often overlook the potential of improving the ability of LLMs to detect and correct misleading content in the input itself. In this paper, we propose a novel three-stage fine-tuning method that enhances the ability of LLMs to detect and correct misleading information in the input, further improving response accuracy and reducing hallucinations. Specifically, the three stages include (1) training LLMs to identify misleading information, (2) training LLMs to correct the misleading information using built-in or external knowledge, and (3) training LLMs to generate accurate answers based on the corrected queries. To evaluate our method, we conducted experiments on three datasets for the hallucination detection task and the question answering (QA) task, as well as two datasets containing misleading information that we constructed. The experimental results demonstrate that our method significantly improves the accuracy and factuality of LLM responses, while also enhancing the ability to detect hallucinations and reducing the generation of hallucinations in the output, particularly when the query contains misleading information. We will publicly release our code upon acceptance.
@@ A Three-Stage Fine-Tuning Method for LLMs
Guocong Li1, Weize Liu1, Yihang Wu1, Ping Wang2, Shuaihan Huang1, Hongxia Xu1,3 Jian Wu1 1Zhejiang University    2Renmin University of China 3AI Research Center, WeDoctor Cloud
@@ 1 Introduction
To evaluate the performance of LLMs in answering queries containing misleading information, we construct two misleading datasets, HaluEval-QAmis and CQAmis, based on the QA subset of the HaluEval dataset and the CQA dataset, respectively. Further, We conduct experiments on the misleading datasets, as well as on the three original datasets for the hallucination detection task and the question answering task. The experimental results show that when input contains misleading information, the performance of the trained model using our method is much higher than that of the original model, demonstrating that our method effectively reduces the impact of misleading information in the input on model performance. Furthermore, regardless of whether the input contains misleading information, our method significantly improves the accuracy and factuality of LLM responses, as well as the ability to detect hallucinations in the responses, while effectively mitigating the hallucinations of the model.
@@ 2.1 Construction of Misleading Query Datasets
The answer error rate measures whether the model generates incorrect answers when responding to generated misleading queries. We input the query into the model $N$ times, record the number of incorrect answers $N_{\text{error}}$, and define the error rate as:
@@
For each query, we randomly generate three misleading queries that satisfy $S_{\text{sim}}$ > 0.8 and $E_{\text{error}}$ > 0.5, and then retain the one with the highest value of $S_{\text{sim}}+E_{\text{error}}$. An example of randomly generated misleading queries is presented in Appendix  C. This ensures that the generated misleading queries are both highly similar to the original queries and contain misleading information, thereby enabling more effective training and evaluation of LLMs’ performance when faced with highly deceptive and subtle misleading queries.
@@ 2.5 Unified Framework and Training
To ensure consistency across the query detection, correction, and answer generation processes, the output of each stage serves as the input for the subsequent stage, ultimately forming a seamless pipeline. Through this three-stage training strategy, the model effectively learns to detect and correct misleading information in queries and generate accurate responses. This approach significantly mitigates the impact and potential harm of misleading information on the model’s performance.
@@ 3 Experiments
Table 1: Performance comparison on HaluEval dataset. In this case, Original refers to the model without any processing, while supervised fine-tuning (SFT) refers to the model that undergoes direct fine-tuning.
@@ 3.1 Setup
  • The HaluEval dataset (Li et al., 2023) includes the QA, Summarization, and Dialogue subsets, is used to test the model’s ability to identify hallucinations and misleading information.
@@
Table 2: Performance comparison on CQA, CQAmis and TruthfulQA dataset.

### v2

@@ Abstract
Large language models (LLMs) exhibit excellent performance in natural language processing (NLP), but remain highly sensitive to the quality of input queries, especially when these queries contain misleading or inaccurate information. Existing methods focus on correcting the output, but they often overlook the potential of improving the ability of LLMs to detect and correct misleading content in the input itself. In this paper, we propose a novel three-stage fine-tuning method that enhances the ability of LLMs to detect and correct misleading information in the input, further improving response accuracy and reducing hallucinations. Specifically, the three stages include (1) training LLMs to identify misleading information, (2) training LLMs to correct the misleading information using built-in or external knowledge, and (3) training LLMs to generate accurate answers based on the corrected queries. To evaluate our method, we conducted experiments on three datasets for the hallucination detection task and the question answering (QA) task, as well as two datasets containing misleading information that we constructed. The experimental results demonstrate that our method significantly improves the accuracy and factuality of LLM responses, while also enhancing the ability to detect hallucinations and reducing the generation of hallucinations in the output, particularly when the query contains misleading information.  [^1] Codes and data are available at: https://github.com/cong03/FMQAA.
@@ A Three-Stage Fine-Tuning Method for LLMs
Guocong Li1, Weize Liu1,  Yihang Wu1, Ping Wang2, Shuaihan Huang1, Hongxia Xu1,3thanks:   Corresponding authors., Jian Wu [^1] footnotemark: 1 1Zhejiang University    2Renmin University of China 3AI Research Center, WeDoctor Cloud guocong_li@163.com    weizeliu1115@gmail.com {einstein,wujian2000}@zju.edu.cn
@@ 1 Introduction
To evaluate the performance of LLMs in answering queries containing misleading information, we construct two misleading datasets, HaluEval-QAmis and CQAmis, based on the QA subset of the HaluEval dataset and the CQA dataset, respectively. Subsequently, we conduct experiments on the misleading datasets, as well as on the three original datasets for the hallucination detection task and the question answering task. The experimental results show that when input contains misleading information, the performance of the trained model using our method is much higher than that of the original model, demonstrating that our method effectively reduces the impact of misleading information in the input on model performance. Furthermore, the model trained using our method also detects that a portion of questions in commonly used standard datasets contain misleading information. Removing these misleading questions can significantly improve the performance of the original model, whereas the model trained with our method consistently maintains stable performance, demonstrating its robustness to misleading information in the input. Regardless of whether the input contains misleading information, our method significantly improves the accuracy and factuality of LLM responses, as well as the ability to detect hallucinations in the responses, while effectively mitigating the hallucinations of the model.
@@ 2.1 Construction of Misleading Query Datasets
The answer error rate measures whether the model generates incorrect answers when responding to generated, misleading queries. We input the query into the model $N$ times, record the number of incorrect answers $N_{\text{error}}$, and define the error rate as:
@@
Specifically, for each query, we randomly generate three misleading queries that satisfy $S_{\text{sim}}$ > 0.8 and $E_{\text{error}}$ > 0.5, and then retain the one with the highest value of $S_{\text{sim}}+E_{\text{error}}$. An example of randomly generated misleading queries is presented in Appendix  C. This ensures that the generated misleading queries are both highly similar to the original queries and contain sufficiently misleading information, thereby enabling more effective training and evaluation of LLMs’ performance when faced with highly deceptive and subtle misleading queries.
@@ 2.5 Unified Framework and Training
To ensure consistency across the query detection, correction, and answer generation processes, we carefully train the model for the next stage based on the model trained in the previous stage, ultimately forming a seamless pipeline. Through this three-stage training strategy, the model effectively learns to detect and correct misleading information in queries and generate accurate responses. This approach significantly mitigates the impact and potential harm of misleading information on the model’s performance.
@@ 3 Experiments
Table 1: Performance (mean $\pm$ standard deviation, %) comparison on HaluEval dataset. In this case, Original refers to the model without any additional processing, while supervised fine-tuning (SFT) refers to the model that undergoes direct fine-tuning.
@@ 3.1 Setup
  • The HaluEval dataset (Li et al., 2023) includes the QA, Summarization, and Dialogue subsets, which are used to test the model’s ability to identify hallucinations and misleading information.
@@
Table 2: Performance (mean $\pm$ standard deviation, %) comparison on CQA, CQAmis and TruthfulQA dataset.
@@ 3.5 Case Study
3.6 Do the questions in standard datasets contain misleading information?
In order to investigate whether some potentially misleading information also exists in commonly used standard datasets (original HaluEval, CQA, and TruthfulQA), we employed the Qwen2-7B model, which had been trained in the first stage, to examine the questions from these standard datasets. Our findings revealed that some questions in the standard datasets contain misleading information. The proportions of questions containing misleading information in each dataset are summarized in Table 6.
Subsequently, we removed the questions identified by the model as containing misleading information from the original datasets to form new datasets (HaluEval-Sumsub, HaluEval-Diasub, HaluEval-QAsub, CQAsub). We then conducted evaluations on these new datasets. The experimental results are shown in Table 7.
We observed that the accuracy of the original Qwen2-7B model’s responses on the datasets from which questions containing misleading information had been removed showed a significant improvement compared to the original datasets. This indicates that the misleading information present in the original datasets can interfere with the model’s ability to generate accurate responses. However, the Qwen2-7B model trained with our method did not show any notable performance changes, demonstrating its robustness against questions that contain misleading information and its reduced susceptibility to such interference.
Dataset Percentage
HaluEval-Sum 3.91%
HaluEval-Dia 14.62%
HaluEval-QA 13.53%
CQA 10.40%
TruthfulQA 6.30%
Table 6: The proportion of questions containing misleading information in the three standard datasets.
Method HaluEval-Sum HaluEval-Dia HaluEval-QA CQA
Original Sub Original Sub Original Sub Original Sub
Qwen2-7B 51.5±0.8 53.6±0.7 70.3±0.4 71.1±0.5 54.5±1.3 57.2±1.1 70.4±0.8 73.1±0.9
Ours 68.1±1.7 68.5±1.4 72.1±1.3 72.3±0.8 68.9±1.6 69.7±0.7 78.8±1.6 79.1±1.1
Table 7: Performance (mean $\pm$ standard deviation, %) comparison before and after addressing misleading information in the dataset. "Original" denotes the original dataset, and "Sub" refers to the dataset subset with misleading information questions removed.